Compute the monthly payments on a 3-year lease for a $28332 car if the annual rate of depreciation is 15% and the leases annual interest rate is 5.7%. Round your answer to the nearest dollar.
Answer #1

Given price of car = $28332

r=57% = 0.57

1st year value = 28332(1+0.57)

=28332(1.57)

=44481.24

with interest

After 1 year value of car = 28332*0.88

=> 24932.16

2nd year value with interest = 24932.16*(1.57)

= 39143.49